OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. Company A charges $25 per month plus $0.22 per minute for cell phone service. Company B charges $35 per month plus $0.12 per minute. How many minutes of usage is needed in one month to make Company B a better deal?

OpenStudy (anonymous):

heygw

OpenStudy (gw2011):

Hi shenita--Here is what I got: Company A: 0.22M+25 Company B: 0.12M+35 These are the two equations that we need where M=minutes 0.22M+25=0.12M+35 0.22M-0.12M=35-25 0.10M=10 M=10/0.10 M=100 minutes I made the two equations equal to each other to determine where the plans are equal. Therefore, at 100 minutes both plans are the same. However, over 100 minutes Company B is the better deal. In other words, at 101 minutes Company B is the better deal.

OpenStudy (anonymous):

okay let me check over

OpenStudy (anonymous):

Y = 0.22X + 25 Y = 0.12X + 35 .22X + 25 = .12X + 35 .1X = 10 X = 100 100 minutes